# Chaos and thermalization in Clifford–Floquet dynamics

### v1

@@ 1. Introduction
Ergodic theory characterizes the long time behavior of deterministic classical dynamical systems and identifies those that behave "chaotically" at long time scales. It also provides quantitative measures of deterministic chaos (for example, the Kolmogorov-Sinai entropy). For a readable introduction to ergodic theory see e.g. [CFS]. There are many examples of classical dynamical systems exhibiting deterministic chaos: special types of billiards [billiards], hyperbolic automorphisms of tori [CFS] and, more generally, Anosov diffeomorphisms and flows [KatokHasselblatt]. Yet another arena for ergodic theory is provided by cellular automata [Pivatoreview].
The theory of deterministic quantum chaos is much less developed. Only the lowest rungs of the classical ergodic hierarchy (ergodicity, weak mixing, mixing) have reasonable quantum counterparts [AlickiFannes]. The existing notions of quantum dynamical entropy [AlickiFannes, CNT] are less motivated than the Kolmogorov-Sinai entropy, and much less is known about them.
@@
To test this, one can begin by looking at an especially tractable type of QCAs: Clifford QCAs. Consider a finite or infinite system of qubits. Its algebra of observables is generated by Pauli matrices $X_{j},Z_{j}$ where the index $j$ labels the qubits. A Clifford QCA is a QCA which maps each Pauli matrix to a monomial in Pauli matrices. One can show that translationally-invariant (TI) Clifford QCAs are essentially equivalent to a special type of classical Cellular Automata (CA) [CQCA1], therefore the methods used to study the latter give insight into the behavior of the former.
The first study of the Floquet dynamics generated by TI Clifford QCAs was undertaken in [CQCA2] and was limited to qubit chains with a single qubit per site. Such TI Clifford QCAs admit a trichotomy into periodic, glider, and fractal types. From the point of view of ergodic theory, fractal QCAs are the most interesting ones. In particular, Ref. [CQCA2] argued that fractal TI Clifford QCAs thermalize arbitrary TI product states. This means that for any local observable $a$ one has
@@
where $\langle\ldots\rangle_{0}$ and $\langle\ldots\rangle_{\infty}$ are the initial product state and the infinite-temperature state, respectively. This property indicates that fractal TI Clifford QCAs are chaotic quantum dynamical systems. [^2] The fractal condition is strictly stronger than the mixing condition. As discussed below, there are many TI Clifford QCAs which are mixing and integrable. These examples illustrate that mixing and chaos should not be identified. Ref. [CQCA2] shows that fractal TI Clifford QCAs also thermalize arbitrary TI Pauli stabilizer states. We are not aware of any other example of a deterministic many-body system, whether classical or quantum, where thermalization was shown for a comparably large class of initial states.
In this note we re-examine the Clifford-Floquet dynamics by making a fuller use of the machinery of classical CAs. First of all, we explain how to generalize most of the results of Ref. [CQCA2] to systems with an arbitrary number of qubits per site and to higher spatial dimensions. The key notion is that of a diffusive CA. It was introduced in [PV1, PV2] and further studied in [ETDS]. Diffusive Clifford QCAs are a generalization of fractal QCAs of Ref. [CQCA2].
@@
Third, we point out a gap in the proof of thermalization of TI product states in [CQCA2]. Specifically, the arguments in Ref. [CQCA2] do not establish that the sequence on the left-hand side of eq. (1) has a limit for an arbitrary fractal QCA. One can prove a weaker statement: a fractal Clifford QCA thermalizes any TI product state for almost all times. This means that the limit exists and is equal to the right-hand side after we exclude a subset of times which is negligible in a suitable sense. We will call this property weak thermalization as opposed to strong thermalization described by eq. (1). We will see that it is much easier to establish weak thermalization for other classes of states as well.
The paper is organized as follows. Section 2 is preparatory. We recall the relation between Clifford QCAs and classical cellular automata (CAs) and discuss Clifford QCAs from the point of view of ergodic theory. We point out that the quantum ergodic hierarchy collapses in this case: ergodic is equivalent to mixing is equivalent to $r$-mixing for all $r>2$. In Section 3 we explain a gap in the proof of Proposition III.2 in [CQCA2] and show how it can be repaired using the results of [ETDS] at the price of weakening the conclusion. We also discuss the growth of operator support under fractal/diffusive Clifford-Floquet dynamics and show that in some cases the support grows only logarithmically with time. In Section 4 we discuss which classes of states are thermalized by diffusive Clifford-Floquet dynamics. This section contains the main new results of the paper. In Section 5 we present numerical evidence that diffusive Clifford QCAs thermalize some states which are not covered by our rigorous results. In Section 6 we discuss our findings.
@@ 2. Clifford QCAs and pseudo-unitary CAs
Let us review the connection between Clifford QCAs and classical CAs following [CQCA1, CQCA2]. Let $N$ be the number of qubits per unit cell and $X_{i}^{\mu},Z_{i}^{\mu}$, $i\in{\mathbb{Z}}^{d}$, $\mu\in\{1,\ldots,N\}$ be the $X$ and $Z$ Pauli matrices acting site-wise. Pauli matrices are generators of the algebra of local observables $\mathscr{A}$}. A basis in this algebra is given by all monomials in Pauli matrices. Such monomials are in one-to-one correspondence with Laurent polynomials in variables $u_{1},\ldots,u_{d}$ with coefficients in ${\mathbb{F}}_{2}^{2N}$, where ${\mathbb{F}}_{2}=\{0,1\}$ is the field with two elements. In what follows we will denote the $d$-tuple $(u_{1},\ldots,u_{d})$ by $u$. If $k\in{\mathbb{Z}}^{d}$, then $u^{k}$ denotes the monomial $u_{1}^{k_{1}}\ldots u_{d}^{k_{d}}$, while $u^{-1}$ denotes $(u_{1}^{-1},\ldots,u_{d}^{-1})$.
@@
Here, it is implicit that the state on the quantum spin chain is the infinite-temperature state $\omega_{\infty}$, while the probability measure for the CA is the normalized Haar measure. [^3] By a state we mean a positive linear functional on the algebra of observables $\mathscr{A}$}. This covers both pure and mixed states. The value of a state $\omega$ on an observable $a$ will be denoted $\omega(a)$. The equivalence of the last three conditions is, of course, well known [Rokhlin, ShirvaniRogers]. Let us show their equivalence to the first three.
@@
Armed with Theorem 1, we can now construct many examples of mixing Clifford QCAs. We use the following result from the theory of linear CAs [ShirvaniRogers].
@@
The condition that $L(u)^{n}$ has no unit eigenvalues for all $n>0$ has the following interpretation on the QCA side: no Pauli monomial behaves periodically with respect to the Clifford-Floquet dynamics. In Ref. [CQCA2] Clifford QCAs which fail the mixing/ergodicity test above were called periodic. Thus, for Clifford QCAs "periodic" is synonymous with non-ergodic or non-mixing.
@@
A more trivial example in 1d is the "shift QCA" which acts by translations to the left or to the right. That is, $L(u)=u^{\pm 1}1_{2N}.$ The mixing property of the shift QCA arises from the fact that every Pauli monomial eventually gets shifted "to infinity". This example illustrates that mixing is a necessary but not a sufficient condition for chaos, and that mixing can coexist with integrability. A similar mechanism operates in the case of the QCA (15): every Pauli monomial eventually decomposes into a configuration of widely separated "gliders" which have a fixed shape and move with a unit speed to the left or to the right [CQCA2].
@@ 3. Strongly and weakly diffusive QCAs
Strongly diffusive CAs were introduced in [PV1], where they are called diffusive CAs.
@@
In general, it seems difficult to check whether a given linear CA or a Clifford QCA is strongly diffusive. There is a weaker property that is easier to check [PV1]. A subset $J$ of the set of natural numbers is said to have density $1$ if $\lim_{n\rightarrow\infty}\frac{1}{n}|J\cap[1,n]|=1$. The complement of a density-1 subset is called a zero-density subset. Intuitively, a density-1 subset comprises almost all natural numbers, while a zero-density subset contains a vanishingly small fraction of naturals. For example, the set of all prime numbers is zero-density, and so is the set of all complete squares.
@@
In Ref. [PV1] weakly diffusive CAs are called diffusive in density.
@@
Ref. [ETDS] showed that weakly diffusive CAs admit an algebraic characterization. [^4] The proof given in [ETDS] is formulated for CAs in one spatial dimension, but it can be easily generalized to arbitrary $d$. We are grateful to B. Hellouin de Menibus for pointing this out.
@@
Non-zero solutions of $L^{n}q=u^{k}q$ are characters of $\Upsilon$ which are $n$-periodic in time up to a spatial translation by $k$ units. In Ref. [ETDS] such characters are called solitons. From the QCA point of view, a soliton corresponds to a Pauli monomial $q$ that is invariant under a combination of some iterate of the QCA $\alpha$ and a spatial translation. Pairs $(n,k)\in{\mathbb{Z}}\times{\mathbb{Z}}^{d}$ label $\mathbb{Z}$}-subgroups of the symmetry ${\mathbb{Z}}\times{\mathbb{Z}}^{d}$ of the problem (time translations times spatial translations). The above theorem shows that a Clifford QCA $\alpha$ is weakly diffusive if and only if no local quantity is preserved by a $\mathbb{Z}$}-symmetry.
Ref. [CQCA2] studied $d=1$, $N=1$ Clifford QCAs. It calls a non-zero solution of the equation $Lq=u^{k}q$ a glider for $M$. Thus saying that $M$ does not have solitons is equivalent to saying that $L^{n}$ does not have gliders for any $n>0$. Ref. [CQCA2] called such Clifford QCAs fractal. Thus, soliton-free Clifford QCAs are a generalization of fractal Clifford QCAs to general $N$ and $d$. Comparing Theorem 3.2 and Theorem 2.2, we also see that every weakly diffusive Clifford QCA is mixing.
@@
Prop. III.2 in [CQCA2] states that if $M$ is fractal, then it strongly thermalizes any translationally-invariant product state. The proof hinges on the claim that $\lim_{n\rightarrow\infty}|L^{n}q|=\infty$ for any non-zero $q\in{\mathbf{M}}$, i.e. on the strong diffusivity of $L(u)$. By Theorem 3.1, the fractal condition is equivalent to weak diffusivity. Thus, the proof of Prop. III.2 can go through as stated only if every weakly diffusive $L$ is strongly diffusive, at least in the case of a single qubit per site and $d=1$.
The argument for strong diffusivity in [CQCA2] involves two steps. First, the sequence $|L^{n}q|$ is shown to be unbounded for any $q\neq 0$, i.e., $\limsup_{n}|L^{n}q|=\infty$. Second, it is argued (by referring to the proof of Prop. III.1) that an infinite sequence of times with a bounded $|L^{n}q|$ forces some power of $L$ to have a glider, leading to a contradiction and the conclusion that $\liminf_{n}|L^{n}q|=\infty$. The first step is valid (the unboundedness of the sequence $|L^{n}q|$ also follows from weak diffusivity). However, the second step appears to be faulty. The quickest way to see this is to consider the following linear CA with a single bit per site:
@@
Hence $|L^{2^{m}}q|=2|q|$ for any $q$ and all sufficiently large $m$ which implies that $L$ is not strongly diffusive. [^5] While this CA is not pseudo-unitary and not even invertible, and does not correspond to a Clifford QCA, the argument for the existence of a glider in [CQCA2] does not use pseudo-unitarity or invertibility.
Lacking a proof of strong diffusivity for a general soliton-free Clifford QCA, one may ask if there are any examples of Clifford QCAs which are provably strongly diffusive. Ref. [ETDS] constructs two examples of invertible 1d CAs which are strongly diffusive. These CAs involve two bits per site and are not pseudo-unitary. But one can turn them into pseudo-unitary CAs by the following "doubling trick". For any $A\in GL(k,{\mathbf{R}})$, we can define a pseudo-unitary element $L\in GL(2k,{\mathbf{R}})$ by letting
@@
It is clear that doubling a strongly diffusive CA gives a strongly diffusive pseudo-unitary CA. Ref. [ETDS] shows that the following two 1d CAs are strongly diffusive:
@@ 4. Some classes of thermalized states
Ref. [CQCA2] studied 1d Clifford QCAs with a single qubit per unit cell and argued that two classes of TI states are strongly thermalized by any soliton-free QCA of this sort: arbitrary TI Pauli stabilizer states and arbitrary TI product states. [^6] Ref. [CQCA2] also shows that glider Clifford QCAs thermalize arbitrary TI Pauli stabilizer states. The proof of the latter relies on strong diffusivity. As we saw in the previous section, there is a gap in the proof of Prop. III.2, hence strong thermalization of arbitrary product states cannot be inferred.
@@
If one of the numbers $|\omega(X_{0})|,|\omega(Y_{0})|,|\omega(Y_{0})|$ is equal to $1$, then $\omega$ is a Pauli stabilizer state and is thermalized in the strong sense (Section III.C of [CQCA2]). Thus, it is sufficient to consider the case when $|\omega(P)|<1$ for any on-site Pauli monomial $P$. Let $\lambda<1$ be the maximal value of $|\omega(P)|$ where $P\in\{X_{0},Y_{0},Z_{0}\}$. By Theorem 3.1, $L$ is weakly diffusive, hence for any non-zero $q\in{\mathbf{M}}$ there is a density-1 $J_{q}\subset{\mathbb{N}}$ such that
@@
As in [CQCA2], one can generalize these results by replacing P-generic product states by states of the form $\omega=\omega_{0}\circ\beta$ where $\omega_{0}$ is a P-generic product state and $\beta$ is an arbitrary Clifford QCA.

### v2

@@ 1. Introduction
Ergodic theory characterizes the long time behavior of deterministic classical dynamical systems and identifies those that behave "chaotically" at long time scales. It also provides quantitative measures of deterministic chaos (for example, the Kolmogorov-Sinai entropy). For a readable introduction to ergodic theory see e.g. [4]. There are many examples of classical dynamical systems exhibiting deterministic chaos: special types of billiards [2], hyperbolic automorphisms of tori [4] and, more generally, Anosov diffeomorphisms and flows [9]. Yet another arena for ergodic theory is provided by cellular automata [13].
The theory of deterministic quantum chaos is much less developed. Only the lowest rungs of the classical ergodic hierarchy (ergodicity, weak mixing, mixing) have reasonable quantum counterparts [1]. While quantum analogs of the Kolmogorov-Sinai entropy have been defined [1, 3], much less is known about them.
@@
To test this, one can begin by looking at an especially tractable type of QCAs: Clifford QCAs. Consider a finite or infinite system of qubits. Its algebra of observables is generated by Pauli matrices $X_{j},Z_{j}$ where the index $j$ labels the qubits. A Clifford QCA is a QCA which maps each Pauli matrix to a monomial in Pauli matrices. One can show that translationally-invariant (TI) Clifford QCAs are essentially equivalent to a special type of classical Cellular Automata (CA) [15], therefore the methods used to study the latter give insight into the behavior of the former.
The first study of the Floquet dynamics generated by TI Clifford QCAs was undertaken in [7] and was limited to qubit chains with a single qubit per site. Such TI Clifford QCAs admit a trichotomy into periodic, glider, and fractal types. From the point of view of ergodic theory, fractal QCAs are the most interesting ones. In particular, Ref. [7] argued that fractal TI Clifford QCAs thermalize arbitrary TI product states. This means that for any local observable $a$ one has
@@
where $\langle\ldots\rangle_{0}$ and $\langle\ldots\rangle_{\infty}$ are the initial product state and the infinite-temperature state, respectively. This property indicates that fractal TI Clifford QCAs are chaotic quantum dynamical systems. [^2] The fractal condition is strictly stronger than the mixing condition. As discussed below, there are many TI Clifford QCAs which are mixing and integrable. These examples illustrate that mixing and chaos should not be identified. Ref. [7] shows that fractal TI Clifford QCAs also thermalize arbitrary TI Pauli stabilizer states. We are not aware of any other example of a deterministic many-body system, whether classical or quantum, where thermalization was shown for a comparably large class of initial states.
In this note we re-examine the Clifford-Floquet dynamics by making a fuller use of the machinery of classical CAs. First of all, we explain how to generalize most of the results of Ref. [7] to systems with an arbitrary number of qubits per site and to higher spatial dimensions. The key notion is that of a diffusive CA. It was introduced in [11, 12] and further studied in [8]. Diffusive Clifford QCAs are a generalization of fractal QCAs of Ref. [7].
@@
Third, we point out a gap in the proof of thermalization of TI product states in [7]. Specifically, the arguments in Ref. [7] do not establish that the sequence on the left-hand side of eq. (1) has a limit for an arbitrary fractal QCA. One can prove a weaker statement: a fractal Clifford QCA thermalizes any TI product state for almost all times. This means that the limit exists and is equal to the right-hand side after we exclude a subset of times which is negligible in a suitable sense. We will call this property weak thermalization as opposed to strong thermalization described by eq. (1). We will see that it is much easier to establish weak thermalization for other classes of states as well.
The paper is organized as follows. Section 2 is preparatory. We recall the relation between Clifford QCAs and classical cellular automata (CAs) and discuss Clifford QCAs from the point of view of ergodic theory. We point out that the quantum ergodic hierarchy collapses in this case: ergodic is equivalent to mixing is equivalent to $r$-mixing for all $r>2$. In Section 3 we explain a gap in the proof of Proposition III.2 in [7] and show how it can be repaired using the results of [8] at the price of weakening the conclusion. We also discuss the growth of operator support under fractal/diffusive Clifford-Floquet dynamics and show that in some cases the support grows only logarithmically with time. In Section 4 we discuss which classes of states are thermalized by diffusive Clifford-Floquet dynamics. This section contains the main new results of the paper. In Section 5 we present numerical evidence that diffusive Clifford QCAs thermalize some states which are not covered by our rigorous results. In Section 6 we discuss our findings.
@@ 2. Clifford QCAs and pseudo-unitary CAs
Let us review the connection between Clifford QCAs and classical CAs following [15, 7]. Let $N$ be the number of qubits per unit cell and $X_{i}^{\mu},Z_{i}^{\mu}$, $i\in{\mathbb{Z}}^{d}$, $\mu\in\{1,\ldots,N\}$ be the $X$ and $Z$ Pauli matrices acting site-wise. Pauli matrices are generators of the algebra of local observables $\mathscr{A}$}. A basis in this algebra is given by all monomials in Pauli matrices. Such monomials are in one-to-one correspondence with Laurent polynomials in variables $u_{1},\ldots,u_{d}$ with coefficients in ${\mathbb{F}}_{2}^{2N}$, where ${\mathbb{F}}_{2}=\{0,1\}$ is the field with two elements. In what follows we will denote the $d$-tuple $(u_{1},\ldots,u_{d})$ by $u$. If $k\in{\mathbb{Z}}^{d}$, then $u^{k}$ denotes the monomial $u_{1}^{k_{1}}\ldots u_{d}^{k_{d}}$, while $u^{-1}$ denotes $(u_{1}^{-1},\ldots,u_{d}^{-1})$.
@@
Here, it is implicit that the state on the quantum spin chain is the infinite-temperature state $\omega_{\infty}$, while the probability measure for the CA is the normalized Haar measure. [^3] By a state we mean a positive linear functional on the algebra of observables $\mathscr{A}$}. This covers both pure and mixed states. The value of a state $\omega$ on an observable $a$ will be denoted $\omega(a)$. The equivalence of the last three conditions is, of course, well known [14, 16]. Let us show their equivalence to the first three.
@@
Armed with Theorem 1, we can now construct many examples of mixing Clifford QCAs. We use the following result from the theory of linear CAs [16].
@@
The condition that $L(u)^{n}$ has no unit eigenvalues for all $n>0$ has the following interpretation on the QCA side: no Pauli monomial behaves periodically with respect to the Clifford-Floquet dynamics. In Ref. [7] Clifford QCAs which fail the mixing/ergodicity test above were called periodic. Thus, for Clifford QCAs "periodic" is synonymous with non-ergodic or non-mixing.
@@
A more trivial example in 1d is the "shift QCA" which acts by translations to the left or to the right. That is, $L(u)=u^{\pm 1}1_{2N}.$ The mixing property of the shift QCA arises from the fact that every Pauli monomial eventually gets shifted "to infinity". This example illustrates that mixing is a necessary but not a sufficient condition for chaos, and that mixing can coexist with integrability. A similar mechanism operates in the case of the QCA (15): every Pauli monomial eventually decomposes into a configuration of widely separated "gliders" which have a fixed shape and move with a unit speed to the left or to the right [7].
@@ 3. Strongly and weakly diffusive QCAs
Strongly diffusive CAs were introduced in [11], where they are called diffusive CAs.
@@
In general, it seems difficult to check whether a given linear CA or a Clifford QCA is strongly diffusive. There is a weaker property that is easier to check [11]. A subset $J$ of the set of natural numbers is said to have density $1$ if $\lim_{n\rightarrow\infty}\frac{1}{n}|J\cap[1,n]|=1$. The complement of a density-1 subset is called a zero-density subset. Intuitively, a density-1 subset comprises almost all natural numbers, while a zero-density subset contains a vanishingly small fraction of naturals. For example, the set of all prime numbers is zero-density, and so is the set of all complete squares.
@@
In Ref. [11] weakly diffusive CAs are called diffusive in density.
@@
Ref. [8] showed that weakly diffusive CAs admit an algebraic characterization. [^4] The proof given in [8] is formulated for CAs in one spatial dimension, but it can be easily generalized to arbitrary $d$. We are grateful to B. Hellouin de Menibus for pointing this out.
@@
Non-zero solutions of $L^{n}q=u^{k}q$ are characters of $\Upsilon$ which are $n$-periodic in time up to a spatial translation by $k$ units. In Ref. [8] such characters are called solitons. From the QCA point of view, a soliton corresponds to a Pauli monomial $q$ that is invariant under a combination of some iterate of the QCA $\alpha$ and a spatial translation. Pairs $(n,k)\in{\mathbb{Z}}\times{\mathbb{Z}}^{d}$ label $\mathbb{Z}$}-subgroups of the symmetry ${\mathbb{Z}}\times{\mathbb{Z}}^{d}$ of the problem (time translations times spatial translations). The above theorem shows that a Clifford QCA $\alpha$ is weakly diffusive if and only if no local quantity is preserved by a $\mathbb{Z}$}-symmetry.
Ref. [7] studied $d=1$, $N=1$ Clifford QCAs. It calls a non-zero solution of the equation $Lq=u^{k}q$ a glider for $M$. Thus saying that $M$ does not have solitons is equivalent to saying that $L^{n}$ does not have gliders for any $n>0$. Ref. [7] called such Clifford QCAs fractal. Thus, soliton-free Clifford QCAs are a generalization of fractal Clifford QCAs to general $N$ and $d$. Comparing Theorem 3.2 and Theorem 2.2, we also see that every weakly diffusive Clifford QCA is mixing.
@@
Prop. III.2 in [7] states that if $M$ is fractal, then it strongly thermalizes any translationally-invariant product state. The proof hinges on the claim that $\lim_{n\rightarrow\infty}|L^{n}q|=\infty$ for any non-zero $q\in{\mathbf{M}}$, i.e. on the strong diffusivity of $L(u)$. By Theorem 3.1, the fractal condition is equivalent to weak diffusivity. Thus, the proof of Prop. III.2 can go through as stated only if every weakly diffusive $L$ is strongly diffusive, at least in the case of a single qubit per site and $d=1$.
The argument for strong diffusivity in [7] involves two steps. First, the sequence $|L^{n}q|$ is shown to be unbounded for any $q\neq 0$, i.e., $\limsup_{n}|L^{n}q|=\infty$. Second, it is argued (by referring to the proof of Prop. III.1) that an infinite sequence of times with a bounded $|L^{n}q|$ forces some power of $L$ to have a glider, leading to a contradiction and the conclusion that $\liminf_{n}|L^{n}q|=\infty$. The first step is valid (the unboundedness of the sequence $|L^{n}q|$ also follows from weak diffusivity). However, the second step appears to be faulty. The quickest way to see this is to consider the following linear CA with a single bit per site:
@@
Hence $|L^{2^{m}}q|=2|q|$ for any $q$ and all sufficiently large $m$ which implies that $L$ is not strongly diffusive. [^5] While this CA is not pseudo-unitary and not even invertible, and does not correspond to a Clifford QCA, the argument for the existence of a glider in [7] does not use pseudo-unitarity or invertibility.
Lacking a proof of strong diffusivity for a general soliton-free Clifford QCA, one may ask if there are any examples of Clifford QCAs which are provably strongly diffusive. Ref. [8] constructs two examples of invertible 1d CAs which are strongly diffusive. These CAs involve two bits per site and are not pseudo-unitary. But one can turn them into pseudo-unitary CAs by the following "doubling trick". For any $A\in GL(k,{\mathbf{R}})$, we can define a pseudo-unitary element $L\in GL(2k,{\mathbf{R}})$ by letting
@@
It is clear that doubling a strongly diffusive CA gives a strongly diffusive pseudo-unitary CA. Ref. [8] shows that the following two 1d CAs are strongly diffusive:
@@ 4. Some classes of thermalized states
Ref. [7] studied 1d Clifford QCAs with a single qubit per unit cell and argued that two classes of TI states are strongly thermalized by any soliton-free QCA of this sort: arbitrary TI Pauli stabilizer states and arbitrary TI product states. [^6] Ref. [7] also shows that glider Clifford QCAs thermalize arbitrary TI Pauli stabilizer states. The proof of the latter relies on strong diffusivity. As we saw in the previous section, there is a gap in the proof of Prop. III.2, hence strong thermalization of arbitrary product states cannot be inferred.
@@
If one of the numbers $|\omega(X_{0})|,|\omega(Y_{0})|,|\omega(Y_{0})|$ is equal to $1$, then $\omega$ is a Pauli stabilizer state and is thermalized in the strong sense (Section III.C of [7]). Thus, it is sufficient to consider the case when $|\omega(P)|<1$ for any on-site Pauli monomial $P$. Let $\lambda<1$ be the maximal value of $|\omega(P)|$ where $P\in\{X_{0},Y_{0},Z_{0}\}$. By Theorem 3.1, $L$ is weakly diffusive, hence for any non-zero $q\in{\mathbf{M}}$ there is a density-1 $J_{q}\subset{\mathbb{N}}$ such that
@@
As in [7], one can generalize these results by replacing P-generic product states by states of the form $\omega=\omega_{0}\circ\beta$ where $\omega_{0}$ is a P-generic product state and $\beta$ is an arbitrary Clifford QCA.